Find the probability of at least 2 girls in 8 births. Assume that male and female births are equally likely and that the births are independent events
Answer by jorel1380(3719)

The probability of at least 2 girls out of 8 births is the same as 1 minus the probability of 8 boys and 7 boys.
8 boys =1/256
7 boys =1/128
8 boys + 7 boys =3/256
Probability of at least 2 girls is 256-3/256=253/256=.98828125..